OpenStudy (anonymous):

please explain how to do 3 5/6+ 2/3 THANKS

OpenStudy (anonymous):

First of all, 3 5/6 is not a number. If you mean 3 and five sixths, it is written 3 + 5/6 or 23 / 6.

OpenStudy (anonymous):

Find a common denominator between 23 / 6 and 2 / 3

OpenStudy (anonymous):

5/6 and 2/3 are the fractions Denominators are 6 and 3 Least common multiples of 6 and 3 is 6 5/6 has denominator 6 We make the denominator of 2/3 to 6 3*2 = 6 We multiply numerator 2 also with 2 2/3 * 2/3 = 4/6 5/6 +4/6 = (5+4)/6 = 9/6 9/6 = 3*3/3*2 = 3/2 = 1.5

OpenStudy (anonymous):

Rita i think (but could be wrong) that he wanted 3 and 5 /6 + 2/3

OpenStudy (anonymous):

mathandphysics is right... could you help? and im a girl-she

OpenStudy (anonymous):

23 / 6 + 2/3 You need to get 2/3 = X / 6 so you have a common denominator. So 2/3 = x / 6, cross multiply to get 3x = 12, so x = 4 So 2/3 = 4/6 Now, 23 / 6 + 4 / 6 = (23+4) / 6 = 27 / 6

OpenStudy (anonymous):

i am C-O-N-F-U-S-E-D!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!